Rewrite each expression as a single trigonometric ratio, and then evaluate the ratio ( show full work )

1a) tan 170• - tan 110• / 1+ tan 170• tan 110•

b) cos 5pie/12 x cos pie over 12 + sin 5pie over 12 x sin pie over 12
1
Expert's answer
2020-10-28T18:47:02-0400

(tan(a) - tan(b))/(1+ tan(a)*tan(b)) = tan(a-b)


(tan(170o) - tan(110o))/(1+ tan(170o)*tan(110o)) = tan(170o-110o) = tan(60o) = sin(60o)/cos(60o) = "\sqrt{3}"


cos(a)*cos(b) + sin(a)*sin(b) = cos(a-b)


cos(5"\pi"/12)*cos("\pi"/12) + sin(5"\pi"/12)*sin("\pi"/12) = cos(5"\pi"/12 - "\pi"/12) = cos(4"\pi"/12) = cos("\pi"/3) = 1/2